# DextrAH-G: Pixels-to-Action Dexterous Arm-Hand Grasping with Geometric Fabrics

### v2

@@ 3.3 Student Depth FGP Training (Policy Distillation)
We use the teacher-student framework and distill our expert to a student policy that can be deployed in the real world using an online version of Dagger [23]. This distillation results in a pixels-to-action policy that uses continuous image input at 15 Hz to perform reactive dynamic grasping in the real world. During distillation, the student $\pi_{depth}(\mathbf{o}_{depth})\rightarrow(\hat{\mathbf{a}},\hat{\mathbf{x}}_{obj})$ receives an observation $\mathbf{o}_{depth}=[\mathbf{o}_{robot},\mathbf{x}_{goal},\mathbf{I}]$, where $\mathbf{I}\in[0.5,1.5]^{160\times 120}$ m is a raw depth image. It produces actions $\hat{\mathbf{a}}\in\mathbb{R}^{11}$ and object position predictions $\hat{\mathbf{x}}_{obj}\in\mathbb{R}^{3}$ ($\hat{\mathbf{x}}_{obj}$ is used by the pick-and-place state machine during real-world deployment). The student is trained with a supervision loss $\mathcal{L}=\mathcal{L}_{action}+\beta\mathcal{L}_{pos}$, where $\mathcal{L}_{action}=||\hat{\mathbf{a}}-\mathbf{a}||_{2}$ and $\mathcal{L}_{pos}=||\hat{\mathbf{x}}_{obj}-\mathbf{x}_{obj}||_{2}$ where $\mathbf{a}$ are actions predicted by the teacher $\pi_{privileged}$ and $\mathbf{x}_{obj}$ are ground-truth object positions from the simulator, and $\beta=0.1$ (see Appendix F). Notably, since the policy fuses depth and proprioceptive signals together, $\hat{\mathbf{x}}_{obj}$ is more accurate through occlusions and facilitates its usage in a state machine. To enable sim2real transfer, we augment the simulated depth readings with noise (see Appendix F.1) to better match the distractors present in the real world.
@@ 4.2 Real-World
Bin Packing Assessment: The bin packing testing protocol quantifies the performance of continuously grasping and transporting a larger variety of objects. Unlike the single-object assessment, this test captures performance within a full application context. Specifically, the robot is charged with continuously grasping over 30 different objects one-at-a-time and transporting them to a bin placed to the side of the robot (see Figures 3, 4). A simple state machine governs the whole process and uses $\hat{\mathbf{x}}_{obj}$ to transition from grasping to transportation (see Appendix G). We propose three primary metrics to quantify grasping performance: 1) consecutive success (CS) - the number of consecutively successful object transports, 2) cycle time - the time required for the robot to grasp an object, transport it, and return to a ready position, and 3) success rate. Across eight different runs, DextrAH-G achieved a CS mean and 95% confidence interval of 6.56 $\pm$ 2.41 objects. This shows that DextrAH-G is capable of transporting several objects in a row before failing (e.g., object falls off the table). Moreover, DextrAH-G also managed a cycle time mean and 95% confidence interval of 10.66 $\pm$ 0.84 seconds, or 5.63 picks-per-minute (PPM) on average. Finally, DextrAH-G successfully grasped and transported all objects with 87% success across 256 attempts. DextrAH-G’s combined celerity and reliability significantly advances the state-of-the art in dexterous robot grasping, bringing real-world utility nearer. For reference, we estimate a human solve rate of 16.53 PPMs for this task based on the Boothroyd-Dewhurst tables with 1.13 s grasp time, 1.5 s repositioning time, and 1 s return time [28]. Under continuous operation, DextrAH-G’s cycle time is already a compelling performance point for practical usage and we expect to improve it in the near future. For additional performance analysis, see Appendix H.
@@ 6 Conclusion
DextrAH-G is a high-performance dexterous grasping policy with depth inputs trained entirely in simulation and deployed in the real world. To achieve this, we combined RL, online distillation, and a vectorized, feature-rich geometric fabric controller. The fabric itself ensures hardware safety and exposes an action space with a strong inductive bias. RL leverages the geometric fabric to learn a privileged FGP for multi-object grasping which is then used to directly train a depth FGP via online distillation. DextrAH-G successfully grasps and transports a large variety of novel objects in the real world, bringing real-world application closer to reality. Critically, over the many hours of testing DextrAH-G (and a variety of ill-behaved FGPs), no hardware was damaged.
@@ B.1 Robot Control via Geometric Fabrics
Geometric fabrics is an artificial dynamical system can be connected with real robots through a torque control law. Industrial, collaborative, and hobbyist robots commonly adopt torque laws that expose a joint position and velocity action space, e.g., joint-level PD control. These controllers track the joint position and velocity targets over time, resulting in closed-loop tracking control. The fabric state above can be used directly as these joint position and velocity targets. Thus, the two dynamical systems become coupled and positively correlated as discussed in [13]. In effect, the real robot’s dynamics are shifted to follow the artificial dynamics closely. Note, in our implementation, the fabric is forward integrated with an approximate second-order Runge-Kutta scheme as in [13] at 60 Hz and we always set the velocity targets to 0.
B.2 Imposing Joint Constraints via Geometric Fabrics
@@ Appendix D Human Motion Retargeting and Fabric PCA Taskmap
An important aspect for a controller is the action space it creates for overlying modules. Within the context of policy learning, the chosen action space can have significant impacts on learning performance and overall behavior. The action space for DextrAH-G enables finger-arm coordination, promotes natural grasping motions, and facilitates RL training for multi-object grasping in simulation at scale. We now focus our attention on part of the action space that involves the fingers. To create an efficient finger action space for reinforcement learning, we find a linear taskmap by running PCA on Allegro finger joint motions derived from retargeting human grasping data. To begin, we leverage a few demonstrations of grasping data from the DexYCB dataset in [47]. This dataset contains 3D point motion traces of human fingertips, joints, and palm of humans throughout object grasping trials. Since the Allegro hand is much bigger than a human hand and only has four fingers, the index, middle, ring, and thumb fingertip points, $\mathbf{x}_{h}\in\mathbb{R}^{12}$, were scaled (scaling factor $\alpha=1.6$) and aligned with the Allegro hand, $\mathbf{x}_{r}\in\mathbb{R}^{12}$ (stacked fingertip points). We aligned the points by creating a palm-fixed coordinate system for the both the human hand and Allegro hand by similarly placing origin and orthonormal axes as in [48]. Points were expressed in their respective palm-fixed coordinate system. We then optimize the following loss, $\mathbb{L}$, that solves for Allegro joint angles, $\mathbf{q}_{r}$, for each data point $\mathbf{x}_{h}$ (in sequence). For the first datapoint, we initialize $\mathbf{q}_{r}$ to zeros. Optimization solves for an unconstrained $\mathbf{q}$, which is then passed through a differentiable saturation function $\mathbf{q}_{r}=\frac{1}{2}(\tanh(\mathbf{q})+1)(\bar{\mathbf{q}}-\underline{\mathbf{q}})+\underline{\mathbf{q}}$, where $\bar{\mathbf{q}}$ and $\underline{\mathbf{q}}$ are the upper and lower joint positional limits of the hand. The loss we optimize via ADAM is
$\mathbb{L}(\mathbf{q})=\gamma\|\mathbf{x}_{r}-\alpha\mathbf{x}_{h}\|^{2}+(1-\gamma)\|\mathbf{x}_{r}-\mathbf{x}_{c}\|^{2}+\lambda||\mathbf{q}_{r}-\mathbf{q}_{reg}\|,$ (4)
@@
where $\mathbf{x}_{c}=[\widetilde{\mathbf{x}}^{T},\widetilde{\mathbf{x}}^{T},\widetilde{\mathbf{x}}^{T},\widetilde{\mathbf{x}}^{T}]^{T}$ stacks $\widetilde{\mathbf{x}}$, where $\widetilde{\mathbf{x}}$ is a single 3D point used for encouraging power or precision grasps. For instance, we place the point on the robot palm to encourage the fingers to fully curl, resulting in a power grasp. We place the point, centrally located among the fingertips to encourage precision grasping. $\mathbf{x}_{r}$ are the stacked Allegro fingertip points by running forward kinematics on $\mathbf{q}_{r}$. $\gamma=1-\frac{i+1}{n}$ is a blend factor that shrinks to 0 over a motion data trace ($i$ is the index of the trace and $n$ is the number of datapoints in a trace). Retargeting for the first data point in a motion trace places all the weight on optimizing the first term in (4). Retaring for the last data point in the motion trace places all the weight in optimizing the second term in (4). This effectively shifts to optimization objective from trying to generate a hand shape that mimics that in the dataset towards one that drives the fingertips towards a power or precision grasp. Finally, $\lambda\in\mathbb{R}^{+}$ is a regularization weight of the retargeted angles towards some fixed, desired angles, $\mathbf{q}_{reg}$. For a precision grip, $\mathbf{q}_{reg}=[0,0,0,0,0,0,0,0,0,0,0,0,1.0,0.75,0,0]$ (an opposed thumb, fingers straight configuration). For a power grip, $\mathbf{q}_{reg}=[0,1,1,1,0,1,1,1,0,1,1,1,1,0.75,0,0]$ (an opposed thumb, fingers curled configuration). Overall, this retargeting process generates a joint-level dataset of Allegro precision and power grasping motions. Please see video for demonstration of retargeted motions.
@@
Given the Allegro grasping motion dataset, we apply PCA to find a rectangular projection matrix with the most dominant eigenvectors. After several applications of PCA, we discover that the five most dominant eigenvectors explain 98% of the variance in the data, which is also consistent with [17]. This rectangular matrix serves as a taskmap within the fabrics framework and exposes a 5-dimensional action space for a fabric-guided policy. Not only did this action space facilitate the learning of high-performance grasping behavior via RL, the grasping behavior itself looked natural as well without any additional reward shaping required. Please see video for fabric behavior that leverages this action space.
@@ E.4 Pose Noise
The policy observes $\widetilde{\mathbf{x}}_{obj}=\mathbf{x}_{obj}+n_{\mathbf{x},uncorr}+n_{\mathbf{x},corr}$ and $\widetilde{\mathbf{q}}_{obj}=\mathbf{q}_{obj}+n_{\mathbf{q},uncorr}+n_{\mathbf{q},corr}$, where $n_{\mathbf{x},uncorr}\sim\mathcal{N}(0,\sigma_{xyz,uncorr})$ and $n_{\mathbf{q},uncorr}\sim\mathcal{N}(0,\sigma_{rpy,uncorr})$ are uncorrelated noise sampled once every simulation timestep, and $n_{\mathbf{x},corr}\sim\mathcal{N}(0,\sigma_{xyz,corr})$ and $n_{\mathbf{q},corr}\sim\mathcal{N}(0,\sigma_{rpy,corr})$ are correlated noise sampled once at the start of each simulation episode and kept the same for the duration of the episode, with $\sigma_{xyz,uncorr}=\sigma_{xyz,corr}=0.02m$ and $\sigma_{rpy,uncorr}=\sigma_{rpy,corr}=0.1rad$. Because poses are represented as quaternions, we convert the roll-pitch-yaw noise to a quaternion then perform quaternion muliplication with the ground truth quaternion.
@@ E.8 Object Datasets
Figure 6: Dataset used in simulation. Objects enumeration starts top left and ends bottom right: e.g. object 15 would be indexed as $[1,1]$ in Python if this was an array data structure.
@@ E.11 Training and Simulation Details
We train our teacher policy with a learning rate of 5e-4, a discount factor $\gamma$ of 0.998, entropy coefficient of 0, and a PPO clipping interval $\epsilon_{clip}$ of 0.2. We also normalize the observations, values, and advantages, and we train the policy with 5 mini-epochs per policy update. Using a horizon length of 16 (number of timesteps between updates for each robot, with all robots running in parallel), and 8192 simulated robots. We use critic MLP of size $[512,512,256,128]$, policy MLP of size $[512,512]$, and policy LSTM of size 1024.
@@ E.13 Comparisons of Grasps
To encourage policy robustness to noise, disturbances, and discrepancies in physics, we apply random force-torque perturbations to the object during grasping, lower contact friction, object pose noise, and significant domain randomization. Like prior works [52, 13], these additions are necessary for strong sim2real transfer for robot manipulation problems. Without it, unnatural and frail strategies emerge as seen in Figures 8,  9, where policies attempt to pick up objects without proper force- and form-closed fingers arrangements. Attempts to grasp objects between the middle and ring fingers only or between the index and ring fingers only were observed. With the above additions, policy strategies become much more robust as seen in Figures 10 and ultimately facilitated DextrAH-G’s high real-world performance.

### v3

@@ 3.3 Student Depth FGP Training (Policy Distillation)
We use the teacher-student framework and distill our expert to a student policy that can be deployed in the real world using an online version of Dagger [23]. This distillation results in a pixels-to-action policy that uses continuous image input at 15 Hz to perform reactive dynamic grasping in the real world. During distillation, the student $\pi_{depth}(\mathbf{o}_{depth})\rightarrow(\hat{\mathbf{a}},\hat{\mathbf{x}}_{obj})$ receives an observation $\mathbf{o}_{depth}=[\mathbf{o}_{robot},\mathbf{x}_{goal},\mathbf{I}]$, where $\mathbf{I}\in[0.5,1.5]^{160\times 120}$ m is a raw depth image. It produces actions $\hat{\mathbf{a}}\in\mathbb{R}^{11}$ and object position predictions $\hat{\mathbf{x}}_{obj}\in\mathbb{R}^{3}$ ($\hat{\mathbf{x}}_{obj}$ is used by the state machine during real-world deployment). The student is trained with a supervision loss $\mathcal{L}=\mathcal{L}_{action}+\beta\mathcal{L}_{pos}$, where $\mathcal{L}_{action}=||\hat{\mathbf{a}}-\mathbf{a}||_{2}$ and $\mathcal{L}_{pos}=||\hat{\mathbf{x}}_{obj}-\mathbf{x}_{obj}||_{2}$ where $\mathbf{a}$ are actions predicted by the teacher $\pi_{privileged}$ and $\mathbf{x}_{obj}$ are ground-truth object positions from the simulator, and $\beta=0.1$ (see Appendix F). Since the policy fuses depth and proprioceptive signals together, $\hat{\mathbf{x}}_{obj}$ is more accurate through occlusions and facilitates its usage in a state machine. We augment the simulated depth readings with noise (see Appendix F.1) to better match the distractors present in the real world.
@@ 4.2 Real-World
Bin Packing Assessment: The bin packing testing protocol quantifies the performance of continuously grasping and transporting a larger variety of objects. Unlike the single-object assessment, this test captures performance within a full application context. Specifically, the robot is charged with continuously grasping over 30 different objects one-at-a-time and transporting them to a bin placed to the side of the robot (see Figures 3, 4). A simple state machine governs the whole process and uses $\hat{\mathbf{x}}_{obj}$ to transition from grasping to transportation (see Appendix G). We propose three primary metrics to quantify grasping performance: 1) consecutive success (CS) - the number of consecutively successful object transports, 2) cycle time - the time required for the robot to grasp an object, transport it, and return to a ready position, and 3) success rate. Across eight different runs, DextrAH-G achieved a CS mean and 95% confidence interval of 6.56 $\pm$ 2.41 objects. This shows that DextrAH-G is capable of transporting several objects in a row before failing (e.g., object falls off the table). Moreover, DextrAH-G also managed a cycle time mean and 95% confidence interval of 10.66 $\pm$ 0.84 seconds, or 5.63 picks-per-minute (PPM) on average. Finally, DextrAH-G successfully grasped and transported all objects with 87% success across 256 attempts. DextrAH-G’s combined celerity and reliability significantly advances the state-of-the art in dexterous robot grasping, bringing real-world utility nearer. For reference, we estimate a human solve rate of 16.53 PPMs for this task based on the Boothroyd-Dewhurst tables with 1.13 s grasp time, 1.5 s repositioning time, and 1 s return time [28]. Under continuous operation, DextrAH-G’s cycle time is already a compelling performance point for practical usage and we expect to improve it in the near future. For additional performance analysis, see Appendix I.
@@ 6 Conclusion
DextrAH-G is a high-performance dexterous grasping policy with depth inputs trained entirely in simulation and deployed in the real world. To achieve this, we combined RL, online distillation, and a vectorized, feature-rich geometric fabric controller. The fabric itself ensures hardware safety and exposes an action space with a strong inductive bias. RL leverages the geometric fabric to learn a privileged FGP for grasping many different objects which is then used to directly train a depth FGP via online distillation. DextrAH-G successfully grasps and transports a large variety of novel objects in the real world, bringing real-world application closer to reality. Critically, over the many hours of testing DextrAH-G (and a variety of ill-behaved FGPs), no hardware was damaged.
Acknowledgments
This work is supported by NVIDIA, DARPA under grant N66001-19-2-4035, NSF Award #1846341, and NSERC Award #526541680.
@@ B.1 Robot Control via Geometric Fabrics
Geometric fabrics is an artificial dynamical system that can be connected with real robots through a torque control law. Industrial, collaborative, and hobbyist robots commonly adopt torque laws that expose a joint position and velocity action space, e.g., joint-level PD control. These controllers track the joint position and velocity targets over time, resulting in closed-loop tracking control. The fabric state above can be used directly as these joint position and velocity targets. Thus, the two dynamical systems become coupled and positively correlated as discussed in [13]. In effect, the real robot’s dynamics are shifted to follow the artificial dynamics closely. Note, in our implementation, the fabric is forward integrated with an approximate second-order Runge-Kutta scheme as in [13] at 60 Hz and we always set the velocity targets to 0.
B.2 Collision Avoidance Details
Next, we provide a more detailed explanation about the formulation of the collision avoidance base metric response per sphere $\mathbf{M}_{b}=\sum_{i}\frac{s_{i}}{\underline{d}_{i}}\hat{\mathbf{n}}_{i}\otimes\hat{\mathbf{n}}_{i}$. The outer product $\hat{\mathbf{n}}_{i}\otimes\hat{\mathbf{n}}_{i}$ creates a matrix with an eigenvector along $\hat{\mathbf{n}}_{i}$. In essence, this matrix will give high priority to actions along $\hat{\mathbf{n}}_{i}$ since that is the direction of importance. Building the complete base metric as a weighted sum over outer products of all $\hat{\mathbf{n}}_{i}$ means that the resulting matrix will have eigenvectors along all $\hat{\mathbf{n}}_{i}$, capturing all directions of importance (directions that point to prospective collision with collision body $i$).
B.3 Imposing Joint Constraints via Geometric Fabrics
@@ B.4 Posture Control
Figure 5: We use a geometric fabric controller with integrated environment and self-collision avoidance. We visualize the geometric fabric’s collision model, which models the robot as a set of spheres and the environment as a set of boxes.
@@ Appendix D Human Motion Retargeting and Fabric PCA Taskmap
An important aspect for a controller is the action space it creates for overlying modules. Within the context of policy learning, the chosen action space can have significant impacts on learning performance and overall behavior. The action space for DextrAH-G enables finger-arm coordination, promotes natural grasping motions, and facilitates RL training for grasping many different objects in simulation at scale. We now focus our attention on part of the action space that involves the fingers. To create an efficient finger action space for reinforcement learning, we find a linear taskmap by running PCA on Allegro finger joint motions derived from retargeting human grasping data. To begin, we leverage a few demonstrations of grasping data from the DexYCB dataset in [47]. This dataset contains 3D point motion traces of human fingertips, joints, and palm of humans throughout object grasping trials. Since the Allegro hand is much bigger than a human hand and only has four fingers, the index, middle, ring, and thumb fingertip points, $\mathbf{x}_{h}\in\mathbb{R}^{12}$, were scaled (scaling factor $\alpha=1.6$) and aligned with the Allegro hand, $\mathbf{x}_{r}\in\mathbb{R}^{12}$ (stacked fingertip points). We aligned the points by creating a palm-fixed coordinate system for both the human hand and Allegro hand by similarly placing origin and orthonormal axes as in [48]. Points were expressed in their respective palm-fixed coordinate system. We then optimize the following loss, $\mathbb{L}$, that solves for Allegro joint angles, $\mathbf{q}_{r}$, for each data point $\mathbf{x}_{h}$ (in sequence). For the first datapoint, we initialize $\mathbf{q}_{r}$ to zeros. Optimization solves for an unconstrained $\mathbf{q}$, which is then passed through a differentiable saturation function $\mathbf{q}_{r}=\frac{1}{2}(\tanh(\mathbf{q})+1)(\bar{\mathbf{q}}-\underline{\mathbf{q}})+\underline{\mathbf{q}}$, where $\bar{\mathbf{q}}$ and $\underline{\mathbf{q}}$ are the upper and lower joint positional limits of the hand. The loss we optimize via ADAM is
$\mathbb{L}(\mathbf{q}_{r})=\gamma\|\mathbf{x}_{r}-\alpha\mathbf{x}_{h}\|^{2}+(1-\gamma)\|\mathbf{x}_{r}-\mathbf{x}_{c}\|^{2}+\lambda||\mathbf{q}_{r}-\mathbf{q}_{reg}\|,$ (4)
@@
where $\mathbf{x}_{c}=[\widetilde{\mathbf{x}}^{T},\widetilde{\mathbf{x}}^{T},\widetilde{\mathbf{x}}^{T},\widetilde{\mathbf{x}}^{T}]^{T}$ stacks $\widetilde{\mathbf{x}}$, where $\widetilde{\mathbf{x}}$ is a single 3D point used for encouraging power or precision grasps. For instance, we place the point on the robot palm to encourage the fingers to fully curl, resulting in a power grasp. We place the point, centrally located among the fingertips to encourage precision grasping. $\mathbf{x}_{r}$ are the stacked Allegro fingertip points by running forward kinematics on $\mathbf{q}_{r}$. $\gamma=1-\frac{i+1}{n}$ is a blend factor that shrinks to 0 over a motion data trace ($i$ is the index of the trace and $n$ is the number of datapoints in a trace). Retargeting for the first data point in a motion trace places all the weight on optimizing the first term in (4). Retargeting for the last data point in the motion trace places all the weight in optimizing the second term in (4). This effectively shifts the optimization objective from trying to generate a hand shape that mimics that in the dataset towards one that drives the fingertips towards a power or precision grasp. Finally, $\lambda\in\mathbb{R}^{+}$ is a regularization weight of the retargeted angles towards some fixed, desired angles, $\mathbf{q}_{reg}$. For a precision grip, $\mathbf{q}_{reg}=[0,0,0,0,0,0,0,0,0,0,0,0,1.0,0.75,0,0]$ (an opposed thumb, fingers straight configuration). For a power grip, $\mathbf{q}_{reg}=[0,1,1,1,0,1,1,1,0,1,1,1,1,0.75,0,0]$ (an opposed thumb, fingers curled configuration). Overall, this retargeting process generates a joint-level dataset of Allegro precision and power grasping motions. Please see video for demonstration of retargeted motions.
@@
Given the Allegro grasping motion dataset, we apply PCA to find a rectangular projection matrix with the most dominant eigenvectors. After several applications of PCA, we discover that the five most dominant eigenvectors explain 98% of the variance in the data, which is also consistent with [17]. This rectangular matrix serves as a taskmap within the fabrics framework and exposes a 5-dimensional action space for a fabric-guided policy. Not only did this action space facilitate the learning of high-performance grasping behavior via RL, the grasping behavior itself looked natural as well without any additional reward shaping required. Please see video for fabric behavior that leverages this action space. Figure 7 shows a visual comparison of example grasps using PCA action spaces of different sizes.
Figure 7: We visualize example grasps in a 2D PCA action space and a 5D PCA action space. The 5D PCA action space enables more diverse grasping behavior while staying on the manifold of human-like grasping motions.
@@ E.4 Pose Noise
The policy observes $\widetilde{\mathbf{x}}_{obj}=\mathbf{x}_{obj}+n_{\mathbf{x},uncorr}+n_{\mathbf{x},corr}$ and $\widetilde{\mathbf{q}}_{obj}=\mathbf{q}_{obj}+n_{\mathbf{q},uncorr}+n_{\mathbf{q},corr}$, where $n_{\mathbf{x},uncorr}\sim\mathcal{N}(0,\sigma_{xyz,uncorr})$ and $n_{\mathbf{q},uncorr}\sim\mathcal{N}(0,\sigma_{rpy,uncorr})$ are uncorrelated noise sampled once every simulation timestep, and $n_{\mathbf{x},corr}\sim\mathcal{N}(0,\sigma_{xyz,corr})$ and $n_{\mathbf{q},corr}\sim\mathcal{N}(0,\sigma_{rpy,corr})$ are correlated noise sampled once at the start of each simulation episode and kept the same for the duration of the episode, with $\sigma_{xyz,uncorr}=\sigma_{xyz,corr}=0.02m$ and $\sigma_{rpy,uncorr}=\sigma_{rpy,corr}=0.1rad$. Because poses are represented as quaternions, we convert the roll-pitch-yaw noise to a quaternion then perform quaternion multiplication with the ground truth quaternion.
@@ E.8 Object Datasets
Figure 8: Dataset used in simulation. Objects enumeration starts top left and ends bottom right: e.g. object 15 would be indexed as $[1,1]$ in Python if this was an array data structure. Circles around the object indicate that they were easiest (green) or hardest (red) for the expert policy to grasp. Further information on per-object success analysis is provided in Appendix F.4
@@ E.11 Training and Simulation Details
We train our teacher policy with a learning rate of 5e-4, a discount factor $\gamma$ of 0.998, entropy coefficient of 0, and a PPO clipping interval $\epsilon_{clip}$ of 0.2. We also normalize the observations, values, and advantages, and we train the policy with 5 mini-epochs per policy update. Using a horizon length of 16 (number of timesteps between updates for each robot, with all robots running in parallel), and 8192 simulated robots. We use a critic MLP of size $[512,512,256,128]$, a policy MLP of size $[512,512]$, and a policy LSTM of size 1024.
@@ E.13 Comparisons of Grasps
To encourage policy robustness to noise, disturbances, and discrepancies in physics, we apply random force-torque perturbations to the object during grasping, lower contact friction, object pose noise, and significant domain randomization. Like prior works [52, 13], these additions are necessary for strong sim2real transfer for robot manipulation problems. Without it, unnatural and frail strategies emerge as seen in Figures 10,  11, where policies attempt to pick up objects without proper force- and form-closed fingers arrangements. Attempts to grasp objects between the middle and ring fingers only or between the index and ring fingers only were observed. With the above additions, policy strategies become much more robust as seen in Figures 12 and ultimately facilitated DextrAH-G’s high real-world performance.